OpenStudy (anonymous):

w=2sqrt2(cos(225)+isin(225)) what is w^4?

jimthompson5910 (jim_thompson5910):

Are you familiar with De Moivre's Theorem?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

@jim_thompson5910 i've got 2sqrt2 (cos(450)+isin(450)) so far

OpenStudy (anonymous):

2sqrt2^4*

jimthompson5910 (jim_thompson5910):

so what is \(\Large \left(2\sqrt{2}\right)^4\) equal to?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

would theta be 900 instead of 450?

jimthompson5910 (jim_thompson5910):

how are you getting 900?

OpenStudy (anonymous):

4 x 225

jimthompson5910 (jim_thompson5910):

oh right, I was looking at the wrong part

jimthompson5910 (jim_thompson5910):

yes, you'll have 64*( cos(900) + i*sin(900) )

jimthompson5910 (jim_thompson5910):

then you'll need to find the angle coterminal to 900 degrees such that this angle is in the interval [0, 360)

OpenStudy (anonymous):

i got 180?

OpenStudy (anonymous):

@jim_thompson5910 is the answer \[-2\sqrt{2}\] ?

jimthompson5910 (jim_thompson5910):

64*( cos(900) + i*sin(900) ) = 64*( cos(180) + i*sin(180) ) 64*( cos(900) + i*sin(900) ) = 64*( -1 + 0 ) 64*( cos(900) + i*sin(900) ) = -64

jimthompson5910 (jim_thompson5910):

so that means \[\Large w^4 = -64\] where \[\Large w = 2\sqrt{2}(\cos(225^{\circ})+i\sin(225^{\circ}))\]

jimthompson5910 (jim_thompson5910):

I'm not sure how you got that result

OpenStudy (anonymous):

thank you